# TaoSR1: The Thinking Model for E-commerce Relevance Search

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Language modelsccs: Computing methodologies Natural language processing
@@ 1. Introduction
Previous approaches primarily relied on BERT-based models, leveraging an encoder-only architecture and bidirectional attention mechanisms to achieve strong text matching capabilities, typically satisfying 80-90% of user search needs. However, the remaining over 10% of long-tail complex queries present a significant challenge, often resulting in inconsistent user experiences. These long-tail queries demand stronger semantic understanding and reasoning capabilities from models. The low-rank problem inherent in BERT’s bidirectional attention limits parameter scaling, thereby constraining the upper bound of BERT model capabilities. Recent research has increasingly explored the application of Large Language Model (LLM) technologies in search systems, including works such as RankLLaMA(10.1145/3626772.3657951, ), LREF(10.1145/3701716.3715246, ) and so on(chen-etal-2025-towards-boosting, ; mehrdad2024large, ). However, most approaches still employ discriminative optimization paradigms or ultimately distill knowledge to BERT models for deployment. As a result, they fall short of effectively addressing queries that require complex reasoning online.
@@
In this paper, we propose a novel framework to optimize and directly deploy a generative LLM for online relevance prediction. However, practical implementation presents significant challenges across three dimensions: (1) Deployment Latency: While Chain-of-Thought (CoT) reasoning enhances model reasoning capabilities and significantly improves the upper bound for solving complex problems, the proportional increase in output tokens correspondingly increases response latency. For relevance models, a single request requires computing relevance between a query and hundreds of candidate items, making real-time online generation of hundreds of CoT sequences computationally prohibitive. (2) Error Accumulation in CoT: Since CoT generates longer intermediate processes, a single hallucination or reasoning error at any intermediate step can propagate and affect final results, leading to incorrect classification outcomes. (3) Discriminative Hallucination: Even with a correct reasoning chain, the model may still produce an incorrect final answer.
@@
To address these challenges, we propose an LLM-based optimization framework comprising three stages: (1) Supervised Fine-Tuning (SFT) with CoT to endow models with reasoning capabilities, improving the upper bound of model discrimination ability; (2) Offline pass@N sampling with Direct Preference Optimization (DPO) to enhance correct answering capability, reducing result errors caused by process errors; (3) Difficulty-based Dynamic Sampling with Group Relative Policy Optimization (GRPO) to further alleviate model discriminative hallucination. Experimental results demonstrate that our trained thinking model significantly outperforms baseline models on a curated offline dataset of challenging queries, exhibiting stronger reasoning and comprehension capabilities. Furthermore, it achieves substantial improvements in online side-by-side (SBS) human evaluations.
@@ 2.1. Relevance Search
Relevance modeling is one of the most fundamental tasks in information retrieval, aiming to model the semantic relevance between queries and items. The field of relevance modeling has experienced extensive historical development. Early approaches focused on feature-based modeling, employing algorithms such as BM25(svore2009machine, ) and TF-IDF(aizawa2003information, ), which relied on manually defined features to address relevance modeling challenges. While these methods proved effective in early information retrieval systems, they lacked cross-scenario generalizability and demanded substantial manual effort. Subsequently, deep learning-based representation modeling methods(shen2014learning, ; 10.1145/3442381.3450129, ) emerged, which utilized deep learning models to encode queries and items as embeddings and further modeled their relevance relationships. However, these approaches lacked the capacity to model broad knowledge and thus possessed limited potential for solving complex cases, with performance remaining suboptimal. The introduction of BERT(devlin2019bert, ) marked a paradigm shift, making fine-tuning of pre-trained language models with Transformer(vaswani2017attention, ) architecture a mainstream approach for relevance modeling(bo2021modeling, ; 10.1145/3534678.3539090, ; 10.1145/3637528.3671559, ). These methods demonstrated the capability to model complex semantic relationships, while their unified task paradigm and end-to-end training architecture significantly reduced the manual effort required for feature engineering. In recent years, decoder-only large language models have exhibited exceptional natural language understanding and generation capabilities, virtually unifying upstream and downstream task paradigms across natural language processing, leading to the emergence of relevance modeling research based on large language models. Notable works including LREF(tang2025lref, ), ProRBP(chen2024towards, ), and research by Mehrdad et al.(mehrdad2024large, ) have employed training methodologies such as supervised fine-tuning (SFT) and Direct Preference Optimization (DPO)(rafailov2023direct, ) to endow large language models with query-item relevance modeling capabilities. However, by adhering to discriminative optimization paradigms or distilling knowledge to smaller BERT-like models for deployment, most prior work fails to effectively address queries that require complex reasoning. In this paper, we address this gap by proposing a framework to optimize a large language model for relevance and deploy it directly in a production online setting.
@@ 2.2. Reasoning Large Language Models
While large language models have demonstrated exceptional capabilities across numerous downstream tasks, conventional large language models still struggle to excel in domains requiring strong reasoning abilities, such as mathematics and programming. To address this limitation, early approaches employed prompt engineering(kojima2022large, ) and supervised fine-tuning(snell2024scaling, ) to guide models in generating step-by-step reasoning chains, thereby improving their reasoning capabilities. However, these methods lacked cross-domain generalizability and failed to fully leverage the rich world knowledge embedded in large language models. The emergence of o1(jaech2024openai, ) and DeepSeek-R1(guo2025deepseek, ) has sparked a surge in using reinforcement learning to enhance the reasoning capabilities of large language models. These reasoning models, including o1 and DeepSeek-R1, employ reinforcement learning algorithms such as PPO(schulman2017proximal, ) or GRPO(shao2024deepseekmath, ) to reward answers sampled from the models themselves, thereby activating the models’ inherent rich world knowledge and improving their performance on challenging reasoning problems in mathematics and programming. Subsequently, additional reinforcement learning algorithms(yu2025dapo, )(zheng2025group, ) and reasoning models(team2025kimi, )(seed2025seed1, ) have been introduced, establishing the use of reinforcement learning algorithms on large language models to enhance reasoning capabilities as a widely adopted technical paradigm. However, current research predominantly focuses on the two domains of mathematics and programming, with limited exploration of reinforcement learning applications in vertical industries and diverse task paradigms. Our work provides an exploratory investigation into the application of reinforcement learning in the e-commerce search domain.
@@ 2.3. Reinforce Learning for Classification
E-commerce search relevance tasks can be formulated as binary or multi-classification problems, where the input consists of query and item features, and the output is the model-predicted relevance label. While numerous studies have explored reinforcement learning for training large language models, most current research focuses primarily on complex generation tasks in mathematics and programming domains, with relatively limited exploration of reinforcement learning applications in generative classification tasks. A particularly relevant work is GenCLS++(he2025gencls++, ), which systematically investigates the performance variations of large language models on generative text classification tasks under different generation and training paradigms. Interestingly, unlike their performance in mathematics and programming domains, in generative classification tasks, models often achieve better performance when they either do not output reasoning paths or place the reasoning process after the output label. Furthermore, conducting reinforcement learning after supervised fine-tuning can yield additional performance improvements. Related research includes the work by Li et al.(li2025think, ), who explored the application of reinforcement learning in multi-modal large language models (MLLMs) for generative classification. Addressing the phenomenon that models perform better without explicit reasoning in generative classification tasks, they further investigated Adaptive-Thinking and achieved promising results. However, current research still lacks deep investigation into the underlying mechanisms behind the performance improvements brought by reinforcement learning. Our work systematically explores the performance gains of SFT, DPO(rafailov2023direct, ), and GRPO(shao2024deepseekmath, ) on generative classification tasks under different training and data configurations, while analyzing the underlying causes of these improvements. This provides an effective reference for applying reinforcement learning in generative classification tasks.
Figure 1. Our proposed TaoSR1 framework comprising three stages: (1) SFT with CoT to endow models with reasoning capabilities; (2) Offline multiple sampling based on a pass@N strategy, combined with DPO, to enhance model generation quality; and (3) Difficulty-based dynamic sampling integrated with GRPO to further mitigate model’s discriminative hallucination problems.:
@@ 3.1.2. RAG-based CoT Generation
Nevertheless, e-commerce relevance annotations are governed by intricate, domain-specific business rules. A single label is insufficient to convey this underlying logic to the model. Simply switching to a generative objective still only leads the model to fit the empirical label distribution of the training data. To address this, we explicitly incorporate a CoT process during SFT to instill domain-specific reasoning capabilities into the model. Our structured CoT framework consists of five sequential steps: (1) Query Understanding, (2) Product Comprehension, (3) Category Matching, (4) Attribute Matching, and (5) Relevance Class Determination. Using the powerful Deepseek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability, ), we synthesized CoT samples based on our existing labeled data. Moreover, business rules in e-commerce relevance are complicated and often customized. For example, retrieving “Mate50 Pro” for the query “Mate50” is considered 3-Related, whereas retrieving “Mate50” for the query “Mate50 Pro” is rated 2-Mismatch. We therefore aim to enable the LLM to more faithfully internalize the business rules in our relevance scenarios.
@@ 3.1.3. SFT with CoT
We experimented with two training paradigms: (1) a “think-then-respond” approach, where the model generates the CoT before the label (¡CoT, label¿), and (2) a “respond-then-think” approach with an output format of ¡label, CoT¿. Counterintuitively, the “think-then-respond” strategy, despite its theoretical advantages, consistently failed to outperform the purely discriminative baseline. Our analysis suggests this is due to error accumulation: while CoT generation elevates the model’s performance ceiling, showing encouraging results on challenging and long-tail cases (e.g., “alternatives to Miu Miu”, where text-only discriminative models often mishandle it by retrieving Miu Miu itself, failing to capture the semantics of “alternatives”), the longer generation process is more susceptible to error accumulation. Minor hallucinations or reasoning deviations in the intermediate steps accumulate and ultimately corrupt the final prediction. In contrast, the “respond-then-think” setting mitigates this issue by predicting the label first, leading to significantly improved metrics. Nevertheless, a performance gap relative to the discriminative baseline still remained. This leads us to hypothesize that CoT-based SFT, while effective at instilling rule-based logic, cannot generalize well across all scenarios. We posit that subsequent reinforcement learning is essential to unlock model potential.
@@ 3.2. Pass@N-based DPO
  • Oracle-Guided Correction for Hard Cases: A key limitation of self-correction is its inability to address cases the model consistently fails to solve. These hard cases with pass@N =0=0, representing the model’s consistent failures, are the most critical targets for optimization. Therefore, for this cohort, we sample from DeepSeek-R1 and designate its correct outputs as the “chosen” responses. These oracle-generated responses are then paired with our model’s own incorrect “rejected” responses to form a supplementary preference dataset $\mathcal{D}^{\prime}_{pass}$. Through this method, approximately 5050% of the original pass@N =0=0 instances were curated into this new set.
@@ 3.3. GRPO with Difficulty-based Sampling
The self-correction component of our DPO strategy is conceptually analogous to GRPO, as both encourage models to explore correct reasoning trajectories through multiple sampling of the model’s own outputs. For the chosen responses in our pass@N=0=0 set, the label-conditioned synthesis process often induced discriminative hallucination, where the generated rationale contradicts the final results. Relative to DPO, GRPO allows for online sampling with larger N, facilitates more diverse exploration to discover more logically coherent reasoning paths. This makes it a natural subsequent step for further optimization. Inspired by DAPO (yu2025dapoopensourcellmreinforcement, ), we propose a difficulty-aware dynamic sampling strategy. A key departure from DAPO is our handling of homogeneous batches: when a sampled set of responses is all correct or all incorrect, we discard the sample entirely rather than resampling. This is because for query-item pairs with definitive relevance relationships, if multiple sampling still fails to produce correct answers, it indicates that the model may lack requisite knowledge. In such cases, the probability of obtaining a correct answer through further sampling becomes extremely low, which would instead slow down overall training efficiency. For groups in which all responses are correct, the within-group relative advantage is zero, leading to vanishing gradients that also degrade training efficiency. Moreover, during online RL we aim to focus learning on difficult instances; therefore, we compute advantages and backpropagate gradients only for groups whose empirical accuracy lies within the range $(0,\gamma)$. The training objective is defined as follows:
@@ 3.4. Cumulative Probability Tiering
Both conventional methods, therefore, share a significant drawback: they necessitate multiple hyperparameters (at least four), which introduces considerable operational complexity in both training and deployment. Furthermore, the manual tuning of these hyperparameters via grid search is computationally expensive and often yields suboptimal configurations. To overcome these limitations, we propose Cumulative Probability Tiering (CumPT), a novel method that extends the standard practice from binary classification—using the positive class probability as the scoring threshold—to the multi-tier setting. CumPT elegantly unifies the process, requiring only a single hyperparameter for online multi-tier assignment, thereby eliminating the complex calibration and thresholding steps. Under the CumPT framework, model training remains unchanged, utilizing the standard cross-entropy objective (Eq.5) to predict probabilities for each of the four relevance classes. This mechanism operates by sequentially accumulating probabilities in descending order of the class tokens (from 4 down to 1) and comparing the sum against a single threshold $\beta_{\text{cum}}$. For instance, if the probability of 4-Excellent $\left(\mathbf{p}_{1}\right)_{\text{id(4)}}$ or cumulative probability $\left(\mathbf{p}_{1}\right)_{\text{id(4)}}+\left(\mathbf{p}_{1}\right)_{\text{id(3)}}$ exceeds $\beta_{\text{cum}}$, the item is assigned to the “Good” tier. If not, we evaluate the cumulative probability $\left(\mathbf{p}_{1}\right)_{\text{id(4)}}+\left(\mathbf{p}_{1}\right)_{\text{id(3)}}+\left(\mathbf{p}_{1}\right)_{\text{id(2)}}$; if this sum exceeds $\beta_{\text{cum}}$, the item is assigned to the “Mid” tier. If this condition also fails, the item is assigned to the “Bad” tier. This procedure is formally outlined in Algorithm 1.

### v4

@@ Abstract.
$c$s: Information systems Language modelsccs: Computing methodologies Natural language processing
@@ 1. Introduction
Previous approaches primarily relied on BERT-based models, leveraging an encoder-only architecture and bidirectional attention mechanisms to achieve strong text matching capabilities, typically satisfying 80-90% of user search needs. However, the remaining over 10% of long-tail complex queries present a significant challenge, often resulting in inconsistent user experiences. These long-tail queries demand stronger semantic understanding and reasoning capabilities from models. The low-rank problem inherent in BERT’s bidirectional attention limits parameter scaling, thereby constraining the upper bound of BERT model capabilities. Recent research has increasingly explored the application of Large Language Model (LLM) technologies in search systems, including works such as RankLLaMA(Ma et al., 2024), LREF(Tang et al., 2025a) and so on(Chen et al., 2025; Mehrdad et al., 2024). However, most approaches still employ discriminative paradigms or distill knowledge to BERT models for deployment. As a result, they fall short of effectively addressing queries that require complex reasoning.
@@
In this paper, we propose a novel framework to optimize and directly deploy a generative LLM for online relevance prediction. However, practical implementation presents significant challenges across three dimensions: (1) Deployment Latency: While Chain-of-Thought (CoT) reasoning enhances model reasoning capabilities and significantly improves the upper bound for solving complex problems, the proportional increase in output tokens correspondingly increases response latency. For relevance models, a single request requires computing relevance between a query and hundreds of candidate items, making real-time online generation computationally prohibitive. (2) Error Accumulation in CoT: Since CoT generates longer intermediate processes, a single hallucination or reasoning error at any intermediate step can propagate and affect final results, leading to incorrect classification outcomes. (3) Discriminative Hallucination: Even with a correct reasoning chain, the model may still produce an incorrect final answer.
@@
To address these challenges, we propose an LLM-based optimization framework comprising three stages: (1) Supervised Fine-Tuning (SFT) with CoT to endow models with reasoning capabilities, improving the upper bound of model discrimination ability; (2) Offline pass@N sampling with Direct Preference Optimization (DPO) to enhance correct answering capability, reducing result errors caused by process errors; (3) Difficulty-based Dynamic Sampling with Group Relative Policy Optimization (GRPO) to further alleviate discriminative hallucination. Experimental results demonstrate that our thinking model significantly outperforms baseline models on a curated dataset of challenging queries, exhibiting stronger reasoning and comprehension capabilities. Furthermore, it achieves substantial improvements in online side-by-side human evaluations.
@@ 2.1. Relevance Search
Relevance modeling is one of the most fundamental tasks in information retrieval, aiming to model the semantic relevance between queries and items. The field of relevance modeling has experienced extensive historical development. Early approaches focused on feature-based modeling, employing algorithms such as BM25(Svore and Burges, 2009) and TF-IDF(Aizawa, 2003), which relied on manually defined features to address relevance modeling challenges. While these methods proved effective in early information retrieval systems, they lacked cross-scenario generalizability and demanded substantial manual effort. Subsequently, deep learning-based representation modeling methods(Shen et al., 2014; Yao et al., 2021) emerged, which utilized deep learning models to encode queries and items as embeddings and further modeled their relevance relationships. However, these approaches lacked the capacity to model broad knowledge and thus possessed limited potential for solving complex cases, with performance remaining suboptimal. The introduction of BERT(Devlin et al., 2019) marked a paradigm shift, making fine-tuning of pre-trained language models with Transformer(Vaswani et al., 2017) architecture a mainstream approach for relevance modeling(Bo et al., 2021; Yao et al., 2022; Lin et al., 2024). These methods demonstrated the capability to model complex semantic relationships, while their unified task paradigm and end-to-end training architecture significantly reduced the manual effort required for feature engineering. In recent years, decoder-only large language models have exhibited exceptional natural language understanding and generation capabilities, virtually unifying upstream and downstream task paradigms across natural language processing, leading to the emergence of relevance modeling research based on large language models. Notable works including LREF(Tang et al., 2025b), ProRBP(Chen et al., 2024), and research by Mehrdad et al.(Mehrdad et al., 2024) have employed training methodologies such as supervised fine-tuning (SFT) and Direct Preference Optimization (DPO)(Rafailov et al., 2023) to endow large language models with query-item relevance modeling capabilities. However, by adhering to discriminative optimization paradigms or distilling knowledge to smaller BERT-like models for deployment, most prior work fails to effectively address queries that require complex reasoning. In this paper, we address this gap by proposing a framework to optimize a large language model for relevance and deploy it directly in a production online setting.
@@ 2.2. Reasoning Large Language Models
While large language models have demonstrated exceptional capabilities across numerous downstream tasks, conventional large language models still struggle to excel in domains requiring strong reasoning abilities, such as mathematics and programming. To address this limitation, early approaches employed prompt engineering(Kojima et al., 2022) and supervised fine-tuning(Snell et al., 2024) to guide models in generating step-by-step reasoning chains, thereby improving their reasoning capabilities. However, these methods lacked cross-domain generalizability and failed to fully leverage the rich world knowledge embedded in large language models. The emergence of o1(Jaech et al., 2024) and DeepSeek-R1(Guo et al., 2025) has sparked a surge in using reinforcement learning to enhance the reasoning capabilities of large language models. These reasoning models, including o1 and DeepSeek-R1, employ reinforcement learning algorithms such as PPO(Schulman et al., 2017) or GRPO(Shao et al., 2024) to reward answers sampled from the models themselves, thereby activating the models’ inherent rich world knowledge and improving their performance on challenging reasoning problems in mathematics and programming. Subsequently, additional reinforcement learning algorithms(Yu et al., 2025b)(Zheng et al., 2025) and reasoning models(Team et al., 2025)(Seed et al., 2025) have been introduced, establishing the use of reinforcement learning algorithms on large language models to enhance reasoning capabilities as a widely adopted technical paradigm. However, current research predominantly focuses on the two domains of mathematics and programming, with limited exploration of reinforcement learning applications in vertical industries and diverse task paradigms. Our work provides an exploratory investigation into the application of reinforcement learning in the e-commerce search domain.
@@ 2.3. Reinforce Learning for Classification
E-commerce search relevance tasks can be formulated as binary or multi-classification problems, where the input consists of query and item features, and the output is the model-predicted relevance label. While numerous studies have explored reinforcement learning for training large language models, most current research focuses primarily on complex generation tasks in mathematics and programming domains, with relatively limited exploration of reinforcement learning applications in generative classification tasks. A particularly relevant work is GenCLS++(He et al., 2025), which systematically investigates the performance variations of large language models on generative text classification tasks under different generation and training paradigms. Interestingly, unlike their performance in mathematics and programming domains, in generative classification tasks, models often achieve better performance when they either do not output reasoning paths or place the reasoning process after the output label. Furthermore, conducting reinforcement learning after supervised fine-tuning can yield additional performance improvements. Related research includes the work by Li et al.(Li et al., 2025), who explored the application of reinforcement learning in multi-modal large language models (MLLMs) for generative classification. Addressing the phenomenon that models perform better without explicit reasoning in generative classification tasks, they further investigated Adaptive-Thinking and achieved promising results. However, current research still lacks deep investigation into the underlying mechanisms behind the performance improvements brought by reinforcement learning. Our work systematically explores the performance gains of SFT, DPO(Rafailov et al., 2023), and GRPO(Shao et al., 2024) on generative classification tasks under different training and data configurations, while analyzing the underlying causes of these improvements. This provides an effective reference for applying reinforcement learning in generative classification tasks.
Figure 1. Our TaoSR1 framework comprising three stages: (1) SFT with CoT to endow models with reasoning capabilities; (2) Offline multiple sampling based on a pass@N strategy, combined with DPO, to enhance model generation quality; and (3) Difficulty-based dynamic sampling integrated with GRPO to further mitigate model’s discriminative hallucination problems.
Our proposed TaoSR1 framework comprising three stages: (1) SFT with CoT to endow models with reasoning capabilities; (2)Offline multiple sampling based on a pass@N strategy, combined with DPO, to enhance model generation quality; and (3) Difficulty-based dynamic sampling integrated with GRPO to further mitigate model’s discriminative hallucination problems.
@@ 3.1.2. RAG-based CoT Generation
Nevertheless, e-commerce relevance annotations are governed by intricate, domain-specific business rules. A single label is insufficient to convey this underlying logic to the model. Simply switching to a generative objective still only leads the model to fit the empirical label distribution of the training data. To address this, we explicitly incorporate a CoT process during SFT to instill domain-specific reasoning capabilities into the model. Our structured CoT framework consists of five sequential steps: (1) Query Understanding, (2) Product Comprehension, (3) Category Matching, (4) Attribute Matching, and (5) Relevance Determination. Using the powerful Deepseek-R1 (DeepSeek-AI et al., 2025), we synthesized CoT samples based on our labeled data. Moreover, business rules in e-commerce relevance are complicated and customized. For example, retrieving “Mate50 Pro” for the query “Mate50” is considered 3-Related, whereas retrieving “Mate50” for the query “Mate50 Pro” is rated 2-Mismatch. We therefore aim to enable LLM to more faithfully internalize the business rules in our scenarios.
@@ 3.1.3. SFT with CoT
We experimented with two training paradigms: (1) a “think-then-respond” approach, where the model generates the CoT before the label (¡CoT, label¿), and (2) a “respond-then-think” approach with an output format of ¡label, CoT¿. Counterintuitively, the “think-then-respond” strategy, despite its theoretical advantages, consistently failed to outperform the discriminative baseline. Our analysis suggests this is due to error accumulation: while CoT generation elevates the model’s performance ceiling, showing encouraging results on challenging long-tail cases (e.g., “alternatives to Miu Miu”, where discriminative models often mishandle it by retrieving Miu Miu itself, failing to capture the semantics of “alternatives”), the longer generation process is more susceptible to error accumulation. Minor hallucinations or reasoning deviations in the intermediate steps accumulate and ultimately corrupt the final prediction. In contrast, the “respond-then-think” setting mitigates this issue by predicting the label first, leading to significantly improved metrics. Nevertheless, a performance gap relative to the discriminative baseline still remained. This leads us to hypothesize that CoT-based SFT, while effective at instilling rule-based logic, cannot generalize well across all scenarios. We posit that subsequent reinforcement learning is essential to unlock model potential.
@@ 3.2. Pass@N-based DPO
  • Oracle-Guided Correction for Hard Cases: A key limitation of self-correction is its inability to address cases the model consistently fails to solve. These hard cases with pass@N =0=0, representing model’s consistent failures, are the most critical targets for optimization. Therefore, for this cohort, we sample from DeepSeek-R1 and designate its correct outputs as “chosen” responses. These oracle-generated responses are then paired with our model’s incorrect “rejected” responses to form supplementary preference dataset $\mathcal{D}^{\prime}_{pass}$. Through this method, approximately 5050% of the original pass@N =0=0 instances were curated into this new set.
@@ 3.3. GRPO with Difficulty-based Sampling
The self-correction component of our DPO strategy is conceptually analogous to GRPO, as both encourage models to explore correct reasoning trajectories through multiple sampling of the model’s own outputs. For the chosen responses in our pass@N=0=0 set, the label-conditioned synthesis process often induced discriminative hallucination, where the generated rationale contradicts the final results. Relative to DPO, GRPO allows for online sampling with larger N, facilitates more diverse exploration to discover more logically coherent reasoning paths. This makes it a natural subsequent step for further optimization. Inspired by DAPO (Yu et al., 2025a), we propose a difficulty-aware dynamic sampling strategy. A key departure from DAPO is our handling of homogeneous batches: when a sampled set of responses is all correct or all incorrect, we discard the sample entirely rather than resampling. This is because for query-item pairs with definitive relevance relationships, if multiple sampling still fails to produce correct answers, it indicates that the model may lack requisite knowledge. In such cases, the probability of obtaining a correct answer through further sampling becomes extremely low, which would instead slow down overall training efficiency. For groups in which all responses are correct, the within-group relative advantage is zero, leading to vanishing gradients that also degrade training efficiency. Moreover, during online RL we aim to focus on difficult instances; therefore, we compute advantages and backpropagate gradients only for groups whose empirical accuracy lies within the range $(0,\gamma)$. The training objective is defined as:
@@ 3.4. Cumulative Probability Tiering
Both conventional methods, therefore, share a significant drawback: they necessitate multiple hyperparameters (at least four), which introduces considerable operational complexity in both training and deployment. Furthermore, the manual tuning of these hyperparameters via grid search is computationally expensive and often yields suboptimal configurations. To overcome these limitations, we propose Cumulative Probability Tiering (CumPT), a novel method that extends the standard practice from binary classification—using the positive class probability as the scoring threshold—to the multi-tier setting. CumPT elegantly unifies the process, requiring only a single hyperparameter for online multi-tier assignment, thereby eliminating the complex calibration and thresholding steps. Under the CumPT framework, model training remains unchanged, utilizing the standard cross-entropy objective (Eq.5) to predict probabilities for each of the four relevance classes. This mechanism operates by sequentially accumulating probabilities in descending order of the class tokens (from 4 down to 1) and comparing the sum against a single threshold $\beta_{\text{cum}}$. For instance, if the probability of 4-Excellent $\left(\mathbf{p}_{1}\right)_{\text{id(4)}}$ or cumulative probability $\left(\mathbf{p}_{1}\right)_{\text{id(4)}}+\left(\mathbf{p}_{1}\right)_{\text{id(3)}}$ exceeds $\beta_{\text{cum}}$, the item is assigned to the “Good” tier. If not, we evaluate the cumulative probability $\left(\mathbf{p}_{1}\right)_{\text{id(4)}}+\left(\mathbf{p}_{1}\right)_{\text{id(3)}}+\left(\mathbf{p}_{1}\right)_{\text{id(2)}}$; if this sum exceeds $\beta_{\text{cum}}$, the item is assigned to the “Mid” tier. If this condition also fails, the item is assigned to “Bad” tier. This procedure is formally outlined in Algorithm 1.